OpenStudy (jmansedg):

Solve: 3 – 4(t + 1) = 7. A. -2 B. C. D. 0

MsBrains (ms-brains):

First subtract 3 from both sides.

OpenStudy (idealist10):

3-4(t+1)=7 4(t+1)=3-7 4(t+1)=-4 t+1=-4/4 t+1=-1 t=-1-1 t=-2 Answer: -2

MsBrains (ms-brains):

Exactly. Good job @Idealist10

OpenStudy (idealist10):

Thank you!